OpenStudy (anonymous):

i really need some help here is the question: -4and 5/6 + 1 and 1/2

OpenStudy (anonymous):

i really need help with this please

OpenStudy (anonymous):

\[-4\frac{ 5 }{ 6 }+1\frac{ 1 }{2 }=\frac{ -(4*6+5) }{ 6}+\frac{ 1*2+1}{ 2}=-\frac{ 29 }{ 6}+\frac{ 3 }{ 2 }=?\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw